You have a particular spring that if you stretch it 10 cm, the potential energy in the spring is 100 Joules. What would be potential energy of this spring be if you were to stretch is twice as far? Give your answer in Joules.
Answer #1

Solution)

1st case, distance, x= 10 cm=0.1 m

Energy, E=100 J

We know,

E=0.5*kx^2=0.5*k*(0.1)^2

2nd case,

E'=0.5*k*(2*0.1)^2=4E

So, E'= 4E=4*100J=400 Joules (Ans)

============